Two waves traveling along a stretched string have the same frequency, but one transports 2.5 times the

power of the other. What is the ratio of the amplitudes of the two waves?


1
Expert's answer
2021-10-24T18:23:28-0400
"E\propto A^2\\\frac{A_2}{A_1}=\sqrt{2.5}=1.58"